what is 15 tbsp converted into drops in pharmacy terms?
1
Expert's answer
2012-10-26T10:49:35-0400
We know that

1 tablespoon = 15 mL

and

20 drops = 1 ml

So,

15 tbsp = 15 * 15 ml = 225 ml = 225 * 20 drops = 4500 drops.